a ball A moving with certain vlocity collides with another body B of same mass at rest.  If the coefficient of restitution is 2/3.  the ratio of velocity of A & B after collision is ?

Grade:

1 Answers

Vikas TU
7 years ago
Momentum is conserved.
Therefore,
mu + 0 = mv1 + mv2
u = v1  + v2 …...............................(1)
 
Inelastic collision,
 
e = 2/3 = (v2 – v1)/(u – 0)
    v2 – v1 = 2u/3 ….........................(2)
 
solve the eqns. (1) and (2) and get the ratio .i.e. v1 / v2.